\vspace{.5cm}

\begin{thebibliography}{AVVV2}

\bibitem{as}
\textsc{M. Abramowitz, and I. Stegun, eds.}:
\textit{Handbook of mathematical functions with formulas,
 graphs and mathematical tables,}  National Bureau of Standards, Dover, New York, 1965.


\bibitem{ael}
\textsc{D. Aharonov, and U. Elias}:
{\it Improved inequalities for trigonometric functions via Dirichlet and Zeta functions},
Math. Ineq. Appl. vol. 16, no. 3 (2013), 851--859.


\bibitem{alz1}
\textsc{H. Alzer}:
\textit{On some inequality of gamma and psi function}, Math. Comp. Vol. 66, 217 (1997), 373--389.


\bibitem{avv1}
\textsc{G.D. Anderson, M.K. Vamanamurthy, and M. Vuorinen}:
\textit{Monotonicity Rules in Calculus}, Amer. Math. Month. Vol. 113, No. 9 (2006), pp. 805--816.

\bibitem{avv2}
\textsc{G.D. Anderson, M.K. Vamanamurthy, and M. Vuorinen}:
\textit{Monotonicity of some functions in calculus}, preprint,\\
https://www.math.auckland.ac.nz/Research/Reports/Series/538.pdf.


\bibitem{avz}
\textsc{G.D. Anderson, M. Vuorinen, and X. Zhang}:
\emph{ Topics in special functions III.}, {\\ http://arxiv.org/abs/1209.1696}.


\bibitem{bar}
\textsc{\'A. Baricz}:
{\it Redheffer Type Inequality for Bessel Functions}, JIPAM. Vol. 8, 1, Art 11, 2007.


\bibitem{bs}
\textsc{M. Becker, and E.L. Strak}:
{\it On a hierarchy of
quolynomial inequalities for $\tan(x)$},  Univ. Beograd. Publ. Elektrotehn Fak Ser
Mat Fiz. No. 602-633 (1978), p. 133--138 .


\bibitem{bk}
\textsc{M. Biernacki, J. Krzy\.z}:
{\it On the monotonicity of certain functionals in the theory of analytic functions},
 Ann. Univ. Mariae. Curie-Sklodowska 2 (1955) 134--145.




\bibitem{czq}
\textsc{C.-P. Chen, J.-W. Zhao, and F. Qi}:
{\it Three inequalities involving hyperbolically trigonometric functions}, RGMIA Res. Rep. Coll. 6 (3) (2003)
437--443. Art. 4.


\bibitem{chen-s}
\textsc{C.-P. Chen, J. S\'andor}:
{\it Inequality chains for Wilker, Huygens and
Lazarevi\'c type inequalities}, {\tt http://ajmaa.org/RGMIA/papers/v15/v15a11.pdf}.


\bibitem{deb}
\textsc{L. Debnath, and C.-J. Zhao}:
{\it New strengthened Jordan's inequality and its applications}, Appl.
Math. Lett., vol. 16, no. 4, 2003, pp. 557--560.



\bibitem{guo}
\textsc{B.-N. Guo, B.-M. Qiao, F. Qi, and W. Li}:
{\it On new proofs of Wilker’s inequalities involving trigonometric
functions}, Math. Ineq. Appl. 6 (2003) 19--22.



\bibitem{jiang}
\textsc{W.-D. Jiang, and H. Yun}:
{\it Sharpening of Jordan’s
Inequality and its Applications},
J. Ineq. Pure and Appl. Math. 7(3) Art. 102, 2006.




\bibitem{hvv}
\textsc{V. Heikkala, M.K. Vamanamurthy, and M. Vuorinen}:
{\it Generalized elliptic integrals},
Comput. Methods and Function Theory
April 2009, Vol. 9, Issue 1, pp 75--109.


\bibitem{ivady}
\textsc{P. Iv\'ady}:
{\it On a beta function inequality},
J. Math. Inequal., 6, 3 (2012) 333--341.

\bibitem{jef}
\textsc{A. Jeffrey}:
{\it Handbook of Mathematical Formulas and Integrals}, 3rd ed., Elsevier Academic Press, 2004.


\bibitem{kvv}
\textsc{R. Kl\'en, M. Visuri, and M. Vuorinen}:
{\it On Jordan type inequalities for hyperbolic functions},
J. Ineq. Appl., vol. 2010, pp. 14.


\bibitem{lili}
\textsc{J.-L. Li, and Y.-L. Li}:
{\it On the Strengthened Jordan's Inequality}, J. Ineq.
Appl. 2007, Art. ID 74328, pp. 8.


\bibitem{lz}
\textsc{L. Li, and J. Zhang}:
{\it A new proof on Redheffer-Williams' inequality},
Far East J. Math. Sci., 56(2011), no.2, 213--217.


\bibitem{mit}
\textsc{D.S. Mitrinovi\'c}:
{\it Analytic Inequalities}, Springer-Verlag, Berlin, 1970.


\bibitem{mor}
\textsc{C. Mortici}:
{\it The natural approach of Wilker-Cusa-Huygens inequalities},
Math. Inequal. Appl., 14, 3 (2011) 535--541.
%%%%%%%%
\bibitem{neua}
\textsc{E. Neuman}:
{\it Refinements and generalizations of certain inequalities involving trigonometric and hyperbolic functions},
Advances in Inequalities and Applications, 1 (2012), No. 1, 1--11.



\bibitem{neu2a}
\textsc{E. Neuman}:
{\it Inequalities Involving Hyperbolic Functions and
Trigonometric Functions}, Bull. Int. Math. Vir. Int.
Vol. 2(2012), 87--92.




\bibitem{neusana}
\textsc{E. Neuman, and J. S\'andor}:
{\it On some inequalities involving trigonometric and hyperbolic functions with
emphasis on the Cusa-Huygens, Wilker, and Huygens inequalities},
 Math. Inequal. Appl. Vol. 13, Number 4 (2010), 715--723.


\bibitem{nejoz1a}
\textsc{E. Neuman, and J. S\'andor}:
{\it Optimal inequalities for hyperbolic and trigonometric functions}, Bull. Math. Anal. Appl. 3(2011),
no. 3, 177--181.



\bibitem{ozban}
\textsc{A. Y. \"Ozban}:
{\it A new refined form of Jordan's inequality and its applications}, Appl. Math.
Lett., vol. 19 (2006), 155--160, 2006.

\bibitem{Pin}
\textsc{I. Pinelis}:
{\it L'Hospital Rules for Monotonicity and the Wilker-Anglesio Inequality},
Amer. Math. Monthly 111 (10) (2004) 905--909.


\bibitem{qi1}
\textsc{F. Qi}:
{\it Extensions and sharpenings of Jordan's and Kober's inequality},
Journal of Mathematics for Technology (in Chinese), 4 (1996), 98--101.


\bibitem{re}
\textsc{R. Redheffer}:
{\it Problem 5642}, Amer. Math. Monthly 76 (1969) 422.


\bibitem{ru}
\textsc{H. Ruskeep\"a\"a}: Mathematica\textsuperscript{\tiny{\textregistered}}
 Navigator. 3rd ed. Academic Press, 2009.



\bibitem{joz111a}
\textsc{J. S\'andor}:
{\it On new refinements of Kober's and Jordan's
trigonometric inequalities}, Notes Number Theory Discrete Math. Vol. 19, 2013, No. 1, 73--83.



\bibitem{joz1a}
\textsc{J. S\'andor}:
{\it Two sharp inequalities for trigonometric and hyperbolic functions}, MIA
15(2012), no.2, 409--413.


\bibitem{joz2}
\textsc{J. S\'andor}:
{\it Sharp Cusa-Huygens and related inequalities}, Notes Number Theory Discrete Math.
Vol. 19, 2013, No. 1, 50--54.


\bibitem{joz3}
\textsc{J. S\'andor, and R. Ol\'ah-G\'al}:
{\it On Cusa-Huygens type trigonometric and
hyperbolic inequalities}, Acta Univ. Sapientiae, Mathematica, 4, 2 (2012) 145--153.

\bibitem{sanben}
\textsc{J. S\'andor, and M. Bencze}:
{\it On Huygens' trigonometric inequality}, RGMIA Res.
Rep. Collection, 8 (2005), No. 3, Art. 14.

%%%%%%%%



%%%%%%%%%%%%%%%%%%%%%%%%%


\bibitem{sandor17b}
\textsc{J. S\'andor}:
{\it On new refinements of Kober's and Jordan's trigonometric inequalities},
Notes Number Th. Discr. Math. 19(2013), no.1, 73--83.

\bibitem{sandor17}
\textsc{J. S\'andor}:
{\it On the concavity of $sin (x)/x$}, Octogon Math. Mag., 13(2005), no.1, 406--407.

\bibitem{sandor17a}
\textsc{J. S\'andor}:
{\it Trigonometric and hyperbolic inequalities}, http://arxiv.org/abs/1105.0859



\bibitem{joz0}
\textsc{J. S\'andor}:
{\it Some integral inequalities}, Elem. Math.
43(1988), 177--180.


\bibitem{joz1}
\textsc{J. S\'andor}:
{\it Sur la fonction Gamma}, P.C.R. Math. Pures Neuch\^atel, Serie I,
21(1989), 4--7.


\bibitem{joz2aa}
\textsc{J. S\'andor}:
{\it On the Gamma function II}, P.C.R. Math. Pures Neuch\^atel, Serie I,
28(1997), 10--12.


\bibitem{joz01}
\textsc{J. S\'andor}:
{\it On the Open problem OQ. 532}, Octogon Math. Mag. 9(2001), No.1B, 569--570.


\bibitem{joz3aa}
\textsc{J. S\'andor}:
{\it On the gamma function III}, P.C.R. Math. Pures Neuch\^atel, Serie 2,
19(2001), 33--40.


\bibitem{sumner}
\textsc{J.S. Sumner, A.A. Jagers, M. Vowe, and J. Anglesio}:
{\it Inequalities involving trigonometric functions}, Amer. Math. Monthly 98 (1991), 264--267.


\bibitem{sz}
\textsc{J. Sun and L. Zhu}:
{\it Six new Redheffer-type inequalities for circular and hyperbolic functions},
Comput. Math. Appl. 56, 2 (2008) 522--529.








\bibitem{wilker}
\textsc{J.B. Wilker}:
{\it Problem E3306}, Amer. Math. Monthly 96 (1989) 55.


%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
\bibitem{wi}
\textsc{J.P. Williams}:
{\it Solution of problem 5642}, Amer. Math. Monthly 76 (1969) 1153--1154.


\bibitem{wu}
\textsc{S.-H. Wu and H.M. Srivastava}:
{\it A weighted and exponential generalization of
Wilker’s inequality and its applications}, Integral Transforms and Spec. Funct.
18 (2007), No. 8, 525--535.


\bibitem{wu2}
\textsc{S.-H. Wu and H.M. Srivastava}:
{\it A further refinement of Wilker's inequality},
Integral Transforms and Spec. Funct.
19 (2008), No, 10, 757--765.


\bibitem{wang}
\textsc{Z.-H.Wang}:
{\it A new Wilker-type inequality},
Journal of Yibing University, vol. 6, pp. 21–22, 2007.


\bibitem{yang}
\textsc{Z.-H. Yang}:
{\it Refinements of a two-sided inequality for trigonometric functions},
J. Math. Inequal., 7, 4 (2013) 601--615



\bibitem{zz}
\textsc{L. Zhang and L. Zhu}:
{\it A new elementary proof of Wilker's inequalities},
Math. Ineq.
Appl. vol. 11 2008, No. 1, 149--151.


\bibitem{z1}
\textsc{L. Zhu}:
{\it Sharpening Redheffer-type inequalities for circular functions},
App. Math. Lett. 22 (2009), 743--748.



\bibitem{z2}
\textsc{L. Zhu}:
{\it Sharpening Jordan's inequality and the Yang Le inequality},
Appl. Math. Lett., 19 (2006), 240–243.


\bibitem{z3}
\textsc{L. Zhu}:
{\it A new simple proof of Wilker's inequality},
Math. Ineq. Appl., vol. 8 (2005)
No. 4, 749–750.


\bibitem{z4}
\textsc{L. Zhu}:
{\it On Wilker-type inequalities},
Math. Ineq. Appl. Vol. 10  (2007) No. 4,
727--731.



\bibitem{sun-z}
\textsc{Z. Sun and L. Zhu}:
{\it On new Wilker-type inequalities},
ISRN Math. Anal.
Vol. 2011, Article ID 681702, 7 pp.





\end{thebibliography}

\documentclass[12pt,a4paper,leqno,verbatim]{amsart}
\newcounter{minutes}
\divide\time by 60
\newcounter{hours}
\multiply\time by 60 \addtocounter{minutes}{-\time}

\usepackage{amssymb}
\usepackage{hyperref}
\usepackage{graphicx}
\usepackage{color}

\date{}
\newfont{\cyrilic}{wncyr10 scaled 1000}
\title[On Jordan's Redheffer's and Wilker's inequality]{On Jordan's Redheffer's and Wilker's inequality}

\author[B.A Bhayo]{Barkat Ali Bhayo}

\author[J. S\'andor]{J\'ozsef S\'andor}

\newcommand{\comment}[1]{}

\swapnumbers
\theoremstyle{plain}

\newtheorem{theorem}[equation]{Theorem}
\newtheorem{lemma}[equation]{Lemma}

\newtheorem{corollary}[equation]{Corollary}

\newtheorem{remark}[equation]{Remark}

\numberwithin{equation}{section}

\begin{document}




\begin{abstract}
In this paper, the authors offer new Jordan, Redheffer and Wilker type
inequalities, along with refinements and converses. Connections with
Euler's gamma function are pointed out, too.

\end{abstract}

\maketitle

%
%
%
%


\section{Introduction}

In the recent years, the refinements of the inequalities of trigonometric functions such as
Adamovi\'c-Mitrinovi\'c, Cusa-Huygens, Jordan inequality, Redheffer inequality,
Becker-Stark inequality, Wilker inequality,
Huygens inequality, and Kober inequality
have been studied extremely by numerous authors,
e.g., see \cite{avv1,avv2,czq,kvv,lili,neua,neu2a,neusana,nejoz1a,joz111a,joz1a,joz2,joz3,sanben}
and the references therein. Motivated by these rapid studies,
in this paper we make a contribution
to the subject by refining the Cusa-Huygens, Jordan and Redheffer inequality, and establish a Wilker type inequality.
In all cases, we
give the upper and lower
bound of $\sin(x)/x$ in terms of simple functions. Meanwhile, we give some Redheffer type inequalities for
trigonometric functions, which refine the existing results in the literature.

For the representation of trigonometric functions in terms of gamma function, we define
for $x,y>0$  the classical \emph{gamma function} $\Gamma$, the {\it digamma function} $\psi$ and the \emph{beta function}  $B(\cdot,\cdot)$ by
   $$\Gamma(x) = \int^\infty_0 e^{-t}t^{x-1}\,dt,\quad
       \psi(x) = \frac{\Gamma'(x)}{\Gamma(x)},\quad
        B(x,y) = \frac{\Gamma(x)\Gamma(y)}{\Gamma(x+y)},$$
respectively. We denote the $trigamma$ function $\psi'$ by $K$, and defined as,
\begin{equation}\label{trigama}
K(x)=\sum_{n=0}^\infty\frac{1}{(x+n)^2},\quad K'(x)=-2\sum_{n=0}^\infty\frac{1}{(x+n)^3}.
\end{equation}
We refer to reader to see \cite{joz0,joz1,joz2aa,joz3aa} for more properties and relations of $K$.
The functions $\Gamma$ and $\psi$ satisfy the recurrence relation
\begin{equation}\label{recgam}
\Gamma(1+z)=z\Gamma(z),\quad \psi(1+z)=z\psi(z).
\end{equation}
The following relation
\begin{equation}\label{recpol}
\psi(1+z)-\psi(z)=\frac{1}{z}
\end{equation}
follows from \eqref{recgam}. Differentiating both sides of \eqref{recpol}, one has
\begin{equation}\label{recK}
K(1+z)-K(z)=-\frac{1}{z^2}.
\end{equation}

The following Euler's reflection formula \cite[6.1.17]{as}
\begin{equation}\label{eulref}
\Gamma(t)\Gamma(1-t)=\frac{\pi}{\sin (\pi t)},\quad 0<t<1,
\end{equation}
can be written as
\begin{equation}\label{sinref}
\frac{x}{\sin(x)}=\Gamma\left(1+\frac{x}{\pi}\right)\Gamma\left(1-\frac{x}{\pi}\right)
=B\left(1+\frac{x}{\pi},1-\frac{x}{\pi}\right),\quad 0<x<\pi.
\end{equation}
The logarithmic differentiation to both sides of \eqref{eulref} gives the following reflection formula,

\begin{equation}\label{eulreftan}
\psi(1-t)-\psi(t)=\frac{\pi}{\tan (\pi t)}.
\end{equation}

Replacing $t$ by $t+1/2$ in \eqref{eulref}, we get
\begin{equation}\label{cosref}
\frac{x}{\cos(x)}=\frac{x}{\pi}\Gamma\left(\frac{1}{2}+\frac{x}{\pi}\right)\Gamma\left(\frac{1}{2}-\frac{x}{\pi}\right)
,\quad 0<x<\frac{\pi}{2}.
\end{equation}


Next we recall the Adamovi\'c-Mitrinovi\'c \cite[p.238]{mit} and Cusa-Huygens \cite{sanben} inqualities
\begin{equation}\label{laztri}
(\cos x)^{1/3}<\frac{\sin x}{x}<\frac{\cos x+2}{3},\quad 0<|x|<\frac{\pi}{2},
\end{equation}
For the refinement of \eqref{laztri}, see \cite{kvv,neua,neusana,joz3,sanben,yang}.

Our first main result refines the inequalities in \eqref{laztri} as follows:

\begin{theorem}\label{thm0803} For $x\in(0,\pi)$, the following inequalities hold,
\begin{equation}\label{inequ0703}
\frac{2-(x/\pi)^3+\cos(x)}{3}<\frac{\sin(x)}{x}<\frac{2-(x/\pi)^4+\cos(x)}{3}.
\end{equation}
\end{theorem}

%

In \cite{wilker}, Wilker asked a question to find the largest constant $c$ such that the inequality
\begin{equation}\label{wilker1003}
a(x)>c\,b(x),\quad 0<|x|<\frac{\pi}{2},
\end{equation}
holds true,
where $$a(x)=\left(\frac{\sin(x)}{x}\right)^2+\frac{\tan(x)}{x}-2,\quad {\rm and}\quad b(x)=x^3\tan(x).$$
The inequality \eqref{wilker1003} is known as the Wilker inequality in the literature. 
Anglesio \cite{sumner}
proved that the ratio $a(x)/b(x)$ is decreasing in $x\in(0,\pi/2)$,
and he answered the question by showing the following two sided
inequality,
\begin{equation}\label{angineq}
2+\frac{8}{45}x^3\tan(x)<\left(\frac{\sin(x)}{x}\right)^2+\frac{\tan(x)}{x}<2+\frac{16}{\pi^4}x^3\tan(x),
\end{equation}
with the best possible constants $8/45$ and $16/\pi^4$. For the new proofs  and refinement of \eqref{angineq},
 see \cite{guo,Pin,wang,wu2,zz,z3}. Thereafter the following Wilker type inequalities,
\begin{eqnarray*}
3+\frac{3}{20}x^3\tan(x)&<&2\,\frac{\sin(x)}{x}+\frac{\tan(x)}{x}
<3+\frac{16}{\pi^4}x^3\tan(x),\,0<x<\frac{\pi}{2},\\
2+\frac{2}{45}x^3\sin(x)&<&\left(\frac{x}{\sin(x)}\right)^2+\frac{x}{\tan(x)}
<2+\left(\frac{2}{\pi}-\frac{16}{\pi^3}\right)x^3\sin(x),\,0<x<\frac{\pi}{2},\\
& &\left(\frac{\sinh(x)}{x}\right)^2+\frac{\tanh(x)}{x}
>2+\frac{8}{45}x^3\tanh(x),\,x>0,\\
& &\left(\frac{x}{\sinh(x)}\right)^2+\frac{x}{\tanh(x)}
<2+\frac{2}{45}x^3\sinh(x),\,x>0,
\end{eqnarray*}
were established by Chen-S\'andor \cite{chen-s}, Wang \cite{wang}, Zhu \cite{z4}, and Sun-Zhu \cite{sun-z},
 respectively.

We establish an other Wilker type inequality by giving the following result.
\begin{theorem}\label{wu-sri} For $x\in(0,\pi)$, we have
$$3+\left(\frac{x}{\pi}\right)^4\frac{x}{\sin(x)}<2\frac{x}{\sin(x)}+\frac{x}{\tan(x)}
<3+\left(\frac{x}{\pi}\right)^3\frac{x}{\sin(x)}.$$
\end{theorem}

The following inequality
\begin{equation}\label{wusri}
2\frac{x}{\sin(x)}+\frac{x}{\tan(x)}>2,\quad 0<|x|<\frac{\pi}{2},
\end{equation}
has been recently established by Wu and Srivastava \cite{wu}, which is sometime known as the second Wilker inequality \cite{neusana}. Inequality \eqref{wusri} can be written as,
$$\frac{\sin(x)}{x}<\frac{1}{2}(\frac{x}{\sin(x)}+\cos(x)),\quad 0<|x|<\frac{\pi}{2}.$$
It was shown in \cite{neusana} that this inequality is weaker than the Cusa-Huygens inequality 
\eqref{laztri}, as follows:
$$\frac{\sin(x)}{x}<\frac{2+\cos(x)}{3}<\frac{1}{2}\left(\frac{x}{\sin(x)}+\cos(x)\right),\quad 0<|x|<\frac{\pi}{2},$$
here second inequality is equivalent to
\begin{equation}\label{ineq1103}
3\,\frac{x}{\sin(x)}+\cos(x)>4,\quad 0<|x|<\frac{\pi}{2}.
\end{equation}
Recently, Mortici \cite{mor} refined this inequality \eqref{ineq1103} as follows:
\begin{equation}\label{neuineq}
3\,\frac{x}{\sin(x)}+\cos(x)>4+\frac{1}{10}x^4+\frac{1}{210}x^6,\quad 0<|x|<\frac{\pi}{2}.
\end{equation}
We establish the following sharp result as a counterpart of \eqref{neuineq}.
\begin{theorem}\label{1003} For $x\in(0,\pi)$, we have
$$2-\frac{x^4}{\pi^4}<3\,\frac{\sin(x)}{x}-\cos(x)<2-\frac{x^3}{\pi^3}.$$
\end{theorem}

The well-know Jordan's inequality \cite{mit} states,
$$\frac{\pi}{2}\leq \frac{\sin(x)}{x},\quad 0<x\leq \frac{\pi}{2},$$
equality with $x=\pi/2$.

In \cite{deb}, Debnath and Zhao refined the Jordan's inequality as below,
\begin{equation}\label{debinequ1}
dz_l(x)=\frac{2}{\pi }+\frac{1}{12\pi}(\pi ^2-4 x^2)\leq\frac{\sin(x)}{x},
  \end{equation}
    \begin{equation}\label{debinequ2}
DZ_l(x)=\frac{2}{\pi }+\frac{1}{\pi^3}(\pi ^2-4 x^2)\leq\frac{\sin(x)}{x},
  \end{equation}
    for $x\in(0,\pi/2)$, equality in both inequalities with $x=\pi/2$.

    The following inequality
    \begin{equation}\label{ozbaninequ1}
 O_l(x)=\frac{2}{\pi }+\frac{1}{\pi^3}(\pi ^2-4 x^2)+\frac{4(\pi-3)}{\pi^3}\left(x-\frac{\pi}{2}\right)^2\leq
\frac{\sin(x)}{x},
  \end{equation}
     for $x\in(0,\pi/2)$ is due to \"Ozban \cite{ozban}, equality with $x=\pi/2$.

    In \cite{z2}, Zhu proved that
    \begin{equation}\label{zinequj}
\frac{\sin(x)}{x}\leq\frac{2}{\pi }+\frac{(\pi -2) \left(\pi ^2-4
   x^2\right)}{\pi^3}=Z_u(x),
\end{equation}
for $x\in(0,\pi/2)$.

For the following inequalities
\begin{equation}\label{doninequ}
J_l(x)=\frac{2}{\pi }+\frac{\pi ^4-16 x^4}{2\pi^5}<\frac{\sin(x)}{x}<\frac{2}{\pi }+\frac{(\pi -2) \left(\pi ^4-16
   x^4\right)}{\pi^5}=J_u(x),
\end{equation}
$x\in(0,\pi/2)$, see \cite{jiang}.

In \cite[Theorem 1.3]{kvv}, Kl\'en et al. proved that
\begin{equation}\label{kvvineq}
1-\frac{x^2}{6}<\frac{\sin(x)}{x}< 1-\frac{2x^2}{3\pi^2},
\end{equation}
for $x\in(-\pi/2,\pi/2)$. The following inequality refines the second inequality in \eqref{kvvineq},
\begin{equation}\label{liineq}
\frac{\sin(x)}{x}\leq \frac{1-x^2/\pi^2}{\sqrt{1+3(x/\pi)^4}},
\end{equation}
for $x\in(0,\pi)$, see \cite{lili}.

Our first main result reads as follows:

\begin{theorem}\label{thm2} For $x\in(0,\pi/2)$, we have
$$C_l(x)<\frac{\sin (x)}{x}<C_u(x),$$
where
$$C_l(x)=\left(1-\frac{x^2}{\pi ^2}\right)^{\pi ^2/6}\quad {\rm and}\quad C_u(x)=\left(1-\frac{x^2}{\pi
   ^2}\right)^{3/2}.$$ 
    \end{theorem}

The right side of the following theorem can not be compared with the corresponding side of Theorem \ref{thm2},
and the second inequality is weaker than the above result.

\begin{theorem}\label{thm1} For $x\in(0,\pi)$, we have
$$D_l(x)<\frac{\sin(x)}{x}<D_u(x),$$
where
$$ D_l(x)=1-\left(\frac{x}{\pi}\right)^2\left(2-\frac{x}{\pi}\right) \quad {\rm and}\quad
D_u(x)=1-\frac{1}{2}\left(\frac{x}{\pi}\right)^2\left(3-\left(\frac{x}{\pi}\right)^2\right).$$
\end{theorem}
It is not difficult to see that Theorem \ref{thm1} refines the inequalities \eqref{kvvineq} and \eqref{liineq}.
Obviously, one can see that $dz_l(x)<DZ_l(x)<O_l(x)$, $J_l(x)<DZ_l(x)$, and
$Z_u(x)<J_u(x)$ for $x\in(0,\pi/2)$. Now it is natural to compare our result with \eqref{ozbaninequ1} and
\eqref{zinequj}. For this purpose
we give the following inequalities by using the Mathematica Software\textsuperscript{\tiny{\textregistered}} \cite{ru},
\begin{eqnarray*}
O_l(x)&<&C_l(x), \quad x\in(0,1.19540),\\
O_l(x)&<&D_l(x), \quad x\in(0,0.92409),\\
C_u(x)&<&Z_u(x), \quad x\in(0,1.09447),\\
D_u(x)&<&Z_u(x), \quad x\in(0,0.95784).
\end{eqnarray*}
We see that our result refines \eqref{ozbaninequ1} and
\eqref{zinequj} in the given interval of $x$.

The following result is the consequence of Theorem \ref{thm1}.
\begin{theorem}\label{gammathm} For $y\in(0,1)$, we have
\begin{enumerate}
\item $\displaystyle B(x,y)<\frac{1}{xy}\frac{x+y}{1+xy},\quad 0<x<1 $,
\item $\displaystyle\frac{1}{xy}\frac{x+y}{1+xy}<B(x,y),\quad 1<x<\infty$.\\
\end{enumerate}
\end{theorem}
It is worth to mention that the part (1) of Theorem \ref{gammathm} recently appeared in \cite{ivady} as a
one of the main results. The proof of the claim is based on
\cite[Lemma 2.5]{ivady}, and the proof of the lemma is invalid.

The following Redheffer inequality \cite{re}
\begin{equation}\label{sinRed}
\frac{\pi^2-x^2}{\pi^2+x^2}\leq \frac{\sin x}{x},\quad 0<x\leq \pi,
\end{equation}
was proved by Williams \cite{wi}.
Chen et al. \cite{czq} obtained the three Redheffer-type inequalities
for $\cos x,\,\cosh x$ and $(\sinh x)/x$. Sun and Zhu \cite{lili,sz} proved the Redheffer-type two-sided inequalities for trigonometric and hyperbolic functions. The inequalities appeared in \cite{lili} were refined in \cite{z2},
and read as follows:

\begin{equation}\label{zhu28a}
\left(\frac{\pi^2-x^2}{\sqrt{\pi^4+3x^4}}\right)^{\pi^2/6}<\frac{\sin(x)}{x}
<\left(\frac{\pi^2-x^2}{\sqrt{\pi^4+3x^4}}\right),\quad 0<x<\pi,
\end{equation}

\begin{equation}\label{zhu28b}
\left(\frac{\pi^2-4x^2}{\sqrt{\pi^4+48x^4}}\right)^{\pi^2/6}<\cos(x)
<\left(\frac{\pi^2-4x^2}{\sqrt{\pi^4+48x^4}}\right)^{3/4},\quad 0<x<\frac{\pi}{2}.
\end{equation}

\begin{equation}\label{zhu28c}
\left(\frac{\sqrt{\pi^4+48x^4}}{\pi^2-4x^2}\right)^{1/2}<\frac{\tan(x)}{x}
<\left(\frac{\sqrt{\pi^4+48x^4}}{\pi^2-4x^2}\right)^{\pi^2/6},\quad 0<x<\frac{\pi}{2}.
\end{equation}
One can see easily that the first inequality in Theorem \ref{thm1} refines the \eqref{sinRed}.
Our following theorem refines \eqref{zhu28a}, as well as Cusa-Huygens inequality,
and also gives a new upper bound for the right side
of \eqref{laztri}.

\begin{theorem}\label{thm3} For $x\in(0,\pi/2)$ one has
\begin{equation}\label{thm3ineq}
\frac{\sin(x)}{x}<\frac{\pi^2-x^2}{\pi^2+\alpha x^2}<\frac{2+\cos (x)}{3}<\frac{\pi^2-x^2}{\pi^2+\beta x^2},
\end{equation}
with best possible constants $\alpha=\pi^2/6-1\approx 0.644934$ and $\beta=1/2$.
\end{theorem}

\begin{theorem}\label{thm4} The function
$$g(t)=\frac{\log((1+t^2)/(1-t^2))}{\log(1/\cos(\pi t/2))}$$
is strictly decreasing from $(0,1)$ on $(\alpha,\beta)$. In particular, for $x\in(0,\pi)$
$$\cos(x/2)^\alpha<\frac{\pi^2-x^2}{\pi^2+x^2}<\cos(x/2)^\beta,$$
with the best possible constants $\alpha=16/\pi^2\approx 1.62114$ and $\beta=1$.
\end{theorem}

In the following theorem we refine the inequalities given in \eqref{zhu28b}.

\begin{theorem}\label{thm1702} The following function
$$h(t)=\frac{1-4t^2}{t^2\cos(\pi t)}-\frac{1}{t^2}$$
is strictly increasing from $(0,1/2)$ onto $((\pi^2-8)/2,16/\pi-4)$. In particular, for $x\in(0,\pi/2)$
$$\frac{\pi^2-4x^2}{\pi^2+\alpha x^2}<\cos(x)<\frac{\pi^2-4x^2}{\pi^2+\beta x^2},$$
with the best possible constants $\alpha=16/\pi-4\approx 1.09296$ and $\beta=(\pi^2-8)/2\approx 0.934802$.
\end{theorem}

The paper is organized into three sections as follows. Section 1,
contains the introduction and the statements of our main results. In Section 2,
we give some lemmas, which will be used
in our proofs.
Section 3 is consists of the proofs of the main results and some corollaries.
\section{Preliminaries}

The following result is sometime called the Monotone l'H\^opital rule, which is due to
Anderson et al. \cite{avv1}.
\begin{lemma}\label{lem0avv}
For $-\infty<a<b<\infty$,
let $f,g:[a,b]\to \mathbb{R}$
be continuous on $[a,b]$, and be differentiable on
$(a,b)$. Let $g^{'}(x)\neq 0$
on $(a,b)$. If $f^{'}(x)/g^{'}(x)$ is increasing
(decreasing) on $(a,b)$, then so are
$$\frac{f(x)-f(a)}{g(x)-g(a)}\quad and \quad \frac{f(x)-f(b)}{g(x)-g(b)}.$$
If $f^{'}(x)/g^{'}(x)$ is strictly monotone,
then the monotonicity in the conclusion
is also strict.
\end{lemma}

\begin{lemma}\label{lem0} For $a\in(0,1)$ and $x>0$, the following inequality holds,
$$\frac{1-a}{x+a}<\psi(1+x)-\psi(x+a).$$
\end{lemma}

\begin{proof} It is well-known that the function $f(x)= x\psi(x),\,x>0$ is strictly convex \cite[Theorem 6]{joz3aa}.
This implies that,
$$f(a\,r+(1-a)s)<a\,f(r)+(1-a)f(s),\quad r,s>0,\,a\in(0,1).$$
Setting $r=1+x$ and $s=x$ in the above inequality, we get
$$(a+x)\psi(a+x)<(a+ax)\psi(1+x)+(1-a)x\psi(x)),$$
now the proof follows easily if we replace $\psi(x)=\psi(1+x)-1/x$.
\end{proof}


\begin{lemma}\label{lem1} For $x\in(1,\infty)$ and $y\in(0,1)$, we have
$$\frac{\Gamma(1+x)\Gamma(1+y)}{\Gamma(1+x+y)}>\frac{1}{1+xy}.$$
\end{lemma}

\begin{proof} Let
$$f_y(x)=\log(\Gamma(1+x))+\log(\Gamma(1+y))-\log(\Gamma(1+x+y))+\log((1+xy)),$$
for $x\in(1,\infty)$ and $y\in(0,1)$, clearly $f_y(1)=0$.
Differentiating $f$ with respect to $x$ and using the formula \eqref{recpol}, we get
\begin{eqnarray*}
f'_y(x)&=&\frac{y}{1+xy}+\psi(1+x)-\psi(1+x+y)\\
&=&\frac{y}{1+xy}-\frac{1}{x+y}+\psi(1+x)-\psi(x+y)\\
&>&\frac{y}{1+xy}-\frac{1}{x+y}+\frac{1-y}{x+y}=\frac{y(1-y)(x-1)}{(x+y)(1+xy)}>0,
\end{eqnarray*}
by Lemma \ref{lem0}.
Thus, $f$ is strictly increasing, and $f_y(x)>f_y(1)=0$, this implies the proof of part (1).
For the proof of part (2), see \cite[(3.2)]{ivady}.
\end{proof}


\begin{remark} \rm
It is easy to see that the convexity of the function $x\mapsto \log(x\Gamma(x)),\,x>0,$
implies the following inequality,
\begin{equation}\label{gru1}
\frac{1}{xy}\left(\frac{x+y}{2}\right)^2<\frac{\Gamma(x)\Gamma(y)}{\Gamma((x+y)/2)^2}.
\end{equation}
Replacing $x$ by $1-x/\pi$ and $y$ by $1-x/\pi$ in \eqref{gru1}, and applying
\eqref{eulref}, we get
\begin{equation}\label{grcon}
\frac{\sin(x)}{x}<\frac{\pi^2-x^2}{\pi^2},\quad 0<x<\pi.
\end{equation}
This improves the  following inequality
$$\frac{\sin(x)}{x}\leq \left(\frac{\pi^2-x^2}{\pi^2+x^2}\right)^{1/2},\quad 0<x\leq \pi,$$
which was proved in \cite{joz01}.
\end{remark}

%

\section{Proof of the main results}

In this section we will give the proofs of the main results
highlighted in the first section, as well as some corollaries are
being established.

\vspace{.3cm}
\noindent{\bf Proof of Theorem \ref{thm0803}.} \rm The proof of the first inequality is trivial.
For the proof of the second and the third inequality, we define
$$f(x)=\frac{2+\cos(x)-3\sin(x)/x}{x^4},$$
for $x\in(0,\pi)$.
We will prove that $f$  is strictly decreasing from $(0, \pi)$ onto $(1/\pi^4, 1/60)$. One has,
$$x^6f(x)=-x^2\sin(x)-7x\cos(x)+15\sin(x)-8x=f_1(x),$$
$$f_1'(x)=5x\sin(x)+8\cos(x)-x^2\cos(x)-8,$$
$$f_2''(x)=-3\sin(x) +3x\cos(x) +x^2\sin (x),\, f_1'''(x)=x(x\cos(x)-\sin(x))<0.$$
Thus $f_1''(x)<f_1''(0)=0$, $f_1'(x)<f_1'(0)=0,$ and $f_1(x)<f_1(0)=0.$ The limiting values can be achieved by l'H\^opital rule.  $\hfill\square$

Similarly, for the proof of the first inequality we will prove that the function
$$g(x)=\frac{2+\cos(x)-3\sin(x)/x}{x^3}$$
is strictly increasing from $(0,\pi)$ onto $(0,1/\pi^3)$.
One has
\begin{eqnarray*}
x^5g'(x)&=& -(x^2)\sin(x) -6x\cos(x)+12 \sin(x)-6x=h(x),\\
h'(x)&= & 4x\sin(x)-x^2\cos(x)+6\cos(x)-6,\\
h''(x)&=& 2x\cos(x) +x^2\sin(x) -2\sin(x), \,h'''(x)= x^2\cos(x).
\end{eqnarray*}
Thus $h'''(x)$ is positive in $x\in(0, \pi/2)$ and negative in $x\in(\pi/2, \pi)$.  Since $h''(\pi/2)=\pi^2/4-2 >0$
 and $h''(\pi)=-2\pi <0$, and $h''(x)$ is strictly increasing (resp. decreasing) in $x\in(0, \pi/2)$
(resp. $x\in(\pi/2,\pi)$). There exists a unique $x_0$ in $(\pi/2, \pi)$ such that $h''(x_0)=0$.
Thus $h''(x)>0,\,x\in (0, x_0)$ and $h''(x)<0,\,x\in(x_0, \pi)$. This implies that $h'(x)$ is strictly increasing
in $x\in(0, x_0)$ and strictly decreasing in $x\in(x_0, \pi)$. As $h'(0)=0$, one has $h'(x_0)>0$.
By  $h'(\pi)= \pi^2-12 <0$, we get that there exists a unique $x_1$ in $(x_0, \pi)$ such that $h'(x_1)=0$.
We get that $h(x)$ is strictly increasing in $(0, x_1)$, and decreasing in $(x_1, \pi)$, with $h(0)=0$ and $h(\pi)=0$.
Thus $h(x)>h(0)=0,\, x\in(0, x_1)$ and $h(x)>h(\pi) =0,\,x\in(x_1, \pi)$. Hence, in all cases, one has  $h(x)>0$.
This completes the proof of the second inequality.

\vspace{.2cm}

It follows from the proof of the second inequality of Theorem \ref{thm0803} that $f<1/60$ and $f_1<0$. This implies the
following result.

\begin{corollary} For $x\in(0,\pi)$, we have
\begin{enumerate}
\item $M_1=\displaystyle\frac{2-x^4/60+\cos(x)}{3}<\frac{\sin(x)}{x}<\frac{8+7\cos(x)}{15-x^2}=M_2,$\\
\item $\displaystyle2\,\frac{x}{\sin(x)}+\frac{x}{\tan(x)}<3+\frac{x^4}{60}\frac{x}{\sin(x)}$,\\
\item $\displaystyle(8/7)\,\frac{x}{\sin(x)}+\frac{x}{\tan(x)}>\frac{15-x^2}{7}$.
\end{enumerate}
\end{corollary}
Recently, the following inequalities appeared in \cite{yang},
\begin{equation}\label{yanginequ}
Y_l=\cos\left(\frac{x}{\sqrt{x}}\right)^a<\frac{\sin(x)}{x}<\cos\left(\frac{x}{\sqrt{x}}\right)^{5/3}=Y_u
\end{equation}
for $x\in(0,\pi/2)$ with $a=\log(2/\pi)/\log(\cos(\sqrt{5}\pi/10))\approx 1.67141$. By using
Mathematica Software\textsuperscript{\tiny{\textregistered}} \cite{ru}, one can see
that
$$Y_l<M_1,\,x\in(0,1.06580)\quad {\rm and}\quad Y_u-M_2\in(0,-0.0009),\, x\in(0,\pi/2).$$

\vspace{.3cm}
\noindent{\bf Proof of Theorem \ref{wu-sri} \& \ref{1003}.} \rm
The proof of both theorems follow immediately from Theorem \ref{thm0803}.

\vspace{.3cm}
\noindent{\bf Proof of Theorem \ref{thm2}.} \rm
Let us consider the application
$$f(x)= \log\left(\frac{x}{\sin(x)}\right)-c \log (1/(1- x^2/\pi^2)),$$
for $x\in(0,\pi/2)$.
A simple computation gives
$$x\sin(x)(\pi^2-x^2)f'(x)= (\sin x-x\cos x)(\pi^2-x^2)-2cx^2\sin(x)=g(x).$$
One has
$$g'(x)/x= (\pi^2-2-4c)\sin( x)+(2-2c)x\cos(x)-x^2\sin(x) = h(x).$$
Finally,
$$h'(x)= (\pi^2-6c)\cos(x)-(4-2c)x\sin(x)-x^2\cos(x).$$
Now, if we select  $c=\pi^2/6$, then, as   $\pi^2-6c=0$ and $4-2c= 4-\pi^2/3>0$, we get $h'(x)<0$, so this finally leads to  $f(x)<0$. Hence, the first inequality follows.

For the proof of the second inequality, let $c=3/2$, then one has
$$g'(x)/x = (\pi^2-8)\sin (x)-x\cos(x) –(x^2)\sin(x)= h(x).$$
Since $h(\pi/2)<0,\, h(\pi/4)>0$, there exists an $x'$ in $(\pi/4, \pi/2)$ such that  $h(x')=0.$ We'll show that
$x'$ is unique. One has $h(x)/\sin (x)= \pi^2-8 -s(x)$, where $s(x)= x^2+ x/\tan(x)$.
Now, $$s'(x)\sin(x)^2 =  2x\sin(x)^2 + \cos (x)\sin( x)-x  =p(x).$$
Here $p'(x)= 4x\sin(x)\cos(x)>0$, which shows that $p(x)>p(0)=0$, so $s'(x)>0$, finally: $s(x)$ is a strictly increasing function. Thus the equation $s(x)= \pi^2-8$  has a single root (which is $x'$), so $h(x)>0$ for $x\in(0, x')$
and $h(x)<0$ for $x\in(x', \pi/2)$. Thus $g$ is increasing, resp. decreasing in the above intervals,
and $g(0)=g(\pi/2)=0$, so $g(x)>0$ for $x\in(0,\pi/2)$. This completes the proof of the second inequality.
$\hfill\square$

\begin{corollary}\label{cor1} For $x\in(0,\pi/2)$ one has
$$\frac{\pi ^2-x^2-\pi ^2 x^2/3}{\pi ^2-x^2}<\frac{x}{\tan(x)}<\frac{\pi ^2-4x^2}{\pi ^2-x^2}.$$
\end{corollary}

\begin{proof} After simplification the derivative of the function
$$f_c(x)= \log\left(\frac{x}{\sin(x)}\right)-c \log (1/(1- x^2/\pi^2)),$$
can be written as
$$f_c'(x)=\frac{1}{x}-{2c\,x}{\pi^2-x^2}-\frac{1}{\tan(x)}.$$
By the proof of Theorem \ref{thm2}, $f_{\pi^2/6}'(x)<0$ and $f_{3/2}'(x)>0$. Clearly,
$f_{\pi^2/6}=0=f_{3/2}$.
Now the proof of the inequalities is obvious.
\end{proof}
The second inequality in Corollary \ref{cor1} improves the first inequality in \eqref{zhu28c}.


\vspace{.3cm}
\noindent{\bf Proof of Theorem \ref{thm1}.} \rm For $t\in(0,1)$, let
$$f(t)=\frac{\pi  t \left(t^3+1\right)-\sin (\pi  t)}{\pi  t^3},\quad
g(t)=\frac{\pi t(2+t^4)-2\sin(\pi t)}{\pi t^3}.$$
By Theorem \ref{thm0803}, we get
\begin{eqnarray*}
f'(x)&=&\frac{3 \pi  t^3+\pi  \left(t^3+1\right)-\pi  \cos (\pi  t)}{\pi  t^3}-\frac{3 \left(\pi  t
   \left(t^3+1\right)-\sin (\pi  t)\right)}{\pi  t^4}\\
&=&\frac{3}{t^3}\left(\frac{\sin(\pi t)}{\pi t}-\frac{2-t^3+\cos(\pi t)}{3}\right)>0,
\end{eqnarray*}
and
\begin{eqnarray*}
g'(x)&=&2t-\frac{2}{t^2}\left(\frac{\cos(\pi t)}{ t}-\frac{\sin(\pi t)}{\pi t^2}\right)+
\frac{4}{t^3}\left(\frac{\sin(\pi t)}{ \pi t}-1\right)\\
&=&\frac{2}{t^3}\left(\frac{\sin(\pi t)}{ \pi t}-\frac{2-t^4+\cos(\pi t)}{3}\right)<0.
\end{eqnarray*}
Thus, the functions $f$ and $g$ are strictly increasing and decreasing in $t\in(0,1)$, respectively. Hence, the proof follows easily if we use the inequalities,
$$f(t)<\lim_{t\to 1}=2,\, g(t)>\lim_{t\to 1}g(t)=3,$$
and replace $t$ by $x/\pi$.$\hfill\square$

\vspace{.3cm}
\noindent{\bf Proof of Theorem \ref{gammathm}}. \rm  Utilizing \eqref{eulref}, the first inequality in Theorem \ref{thm1} is equivalent to
$$\frac{1+x/\pi-x/\pi}{\Gamma(1+x/\pi)\Gamma(1-x/\pi)}>\frac{\left(1-x/\pi\right)\left(1+x/\pi-
(x/\pi)^2\right)}{\Gamma(1+x/\pi-x/\pi)}.$$ Replacing $x$ by $x/\pi$
and $y$ by $1-x/\pi$ we get (1). Similarly, the second inequality in
Theorem \ref{thm1} can be written as,
$$\frac{1+x/\pi+1-x/\pi}{(1-(x/\pi)^2)(2+(x/\pi)^2)} < \frac{\Gamma(1+x/\pi)\Gamma(1-x/\pi)}{\Gamma(1+x/\pi-x/\pi)}.$$
If we replace $x$ by $1+x/\pi$ and $y$ by $1-x/\pi$ then we get the proof of part (2) for $1<x<2$. The rest of proof follows from Lemma \ref{lem1} .$\hfill\square$

%


\vspace{.3cm}
\noindent{\bf Proof of Theorem \ref{thm3}.} \rm Let $f(t)=f_1(t)/f_2(t),\,t\in(0,1/2)$, where
$$f_1(t)=1-3t^2-\cos(\pi t)\quad {\rm and}\quad f_2(t)=t^2(2+\cos(\pi t)).$$
A simple calculation gives
$$f'_1(t)=-6t-\pi\sin(\pi t)<0, \quad {\rm and}\quad f'_2(t)=t(4+2\cos(\pi t)-\pi t\sin(\pi t))>0.$$
Thus, $f$ is the product of two positive strictly decreasing functions,
this implies that $f$ is strictly decreasing in $t\in(0,1/2)$.
Applying l'H\^opital rule, we get
$\lim_{t\to 1/2}f(t)=1/2<f(t)<\alpha=\lim_{t\to 0}f(t)=\pi^2/6-1$. Here the first inequality implies
$$\frac{3(1-t^2)}{2+\cos(\pi t)}-1<\alpha t^2,$$
which is equivalent to
$$\frac{2+\cos(\pi t)}{3}>\frac{1-t^2}{1+\alpha t^2}.$$
Letting $\pi t=x\in(0,\pi/2)$, we get the second inequality of \eqref{thm3ineq},
and the third inequality of \eqref{thm3ineq} follows similarly. For the proof of the first inequality,
see \cite[(2.5)]{ael}. $\hfill\square$

\begin{corollary}\label{coro2} For $x\in(0,1)$, we have

\begin{equation}\label{stark}
\frac{4}{\pi}\frac{t}{1-t^2}<\tan\left(\frac{\pi t}{2}\right)<\frac{\pi}{2}\frac{t}{1-t^2}.
\end{equation}
\end{corollary}

\begin{proof} Let $f(t)=t/(\tan(\pi t/2)(1-t^2)),\,t\in(0,1/2)$. We get
$$f'(t)=\frac{(1+t^2)\sin(\pi t)-\pi t(1-t^2)}{2(1-t^2)^2\sin(\pi t/2)^2},$$
which is positive by \eqref{sinRed}. By l'H\^opital rule, we get
$\lim_{t\to 0}f(t)=2/\pi <f(t)<\pi/4=\lim_{t\to 1/2}f(t)$. This completes the proof.
\end{proof}
For $0<t<1$, letting $x=(\pi t)/2$ in \eqref{stark}, we get
\begin{equation}\label{starkineq}
\frac{8}{\pi^2-4x^2}<\frac{\tan(x)}{x}<\frac{\pi^2}{\pi^2-4x^2},
\end{equation}
which is so-called Becker-Stark inequality \cite{bs}. Thus, Corollary \ref{coro2} gives a simple proof of Becker-Stark inequality. It is easy to see that the second inequality in Corollary \ref{cor1} improves the first inequality
in \eqref{starkineq}.


\vspace{.3cm}
\noindent{\bf Proof of Theorem \ref{thm4}}. \rm Write $g(t)=g_1(t)/g_2(t),\ 0<t<1$, where
$$g_1(t)=\log\left(\frac{1+t^2}{1-t^2}\right),\,g_2(t)=\log\left(\frac{1}{\cos(\pi t/2)}\right).$$
We get,
$$\frac{g'_1(t)}{g'_2(t)}=\frac{8}{\pi}\frac{t}{(1-t^4)\tan(\pi t/2)}=\frac{8}{\pi}g_3(t).$$ One has,
$$g'_3(t)=\frac{\sin(\pi t)(1+3t^4)-\pi t(1-t^4)}{2\sin(\pi t/2)^2(1-t^4)^2},$$
which is negative by \eqref{zhu28a}.
Clearly $g_1(0)=0=g_2(0),$ thus $g$ is strictly decreasing by Lemma \ref{lem0avv}.
Using l'H\^opital rule, we get
$\lim_{t\to 0}g(t)=16/\pi^2 >g(t)>1=\lim_{t\to 1}g(t)$. Replacing $\pi t$ by $x$, we get the desired inequalities.$\hfill\square$

\begin{corollary}\label{coro3} For $x\in(0,\pi)$, one has
$$\left(\frac{\pi^2-x^2}{\pi^2+x^2}\right)^{4/(3\beta)}<\cos\left(\frac{x}{2}\right)^{4/3}<\frac{\sin(x)}{x}<
\left(\frac{\pi^2-x^2}{\pi^2+x^2}\right)^{4/(3\alpha)},$$
where $\alpha$ and $\beta$ are as in Theorem \ref{thm4}.
\end{corollary}

\begin{proof} The proof of the first inequality follows from Theorem \ref{thm4}, and the second inequality is
also well known \cite{neusana}. The third inequality is just \eqref{zhu28a}.
\end{proof}

\begin{theorem} For $x\in(0,\pi/2)$, we have
\begin{equation}\label{sinmax}
\frac{2}{\pi}+\frac{\pi-2}{\pi}\cos(x) < \frac{\sin(x)}{x},
\end{equation}
\begin{equation}\label{sinold}
\left(\frac{1+\cos}{2}\right)^{2/3} < \frac{\sin(x)}{x}< \frac{2\cdot2^{2/3}}{\pi}\left(\frac{1+\cos}{2}\right)^{2/3}
<\frac{4}{\pi}\frac{1+\cos(x)}{2}.
\end{equation}
\end{theorem}

\begin{proof}
The inequality \eqref{sinmax} may be rewritten as
$$f(x)= \pi\sin(x)-(\pi-2)x\cos(x)- 2x>0,$$
for $x\in(0,\pi/2)$.
Clearly, $f(0)=f(\pi/2)=0.$
On the other hand,

$$f(x)= 2(\cos (x)-1) +(\pi-2)x\sin(x) = -4\sin (x/2)^2 +2(\pi-2)x\sin (x/2)\cos (x/2).$$
 This implies that
$$f'(x)/(4\sin(x/2)\cos(x/2))= (\pi-2)/2-\tan( x/2) =g(x) $$
for $x>0$.
As $\tan (x/2)$ is strictly increasing, there is a unique $x'$ in $(0,\pi/2)$ such that $g(x')=0$. Also,
$g(x)>0$  for $x\in(0,x')$ and $g(x)<0$  for $x\in(x', \pi/2)$; i.e.  $x'$ is a maximum point of $f(x)$.
This gives $f(x)>f(0)=0$  for $x\in (0,x')$ and $f(x)>f(\pi/2)=0$  for $x \in (x',\pi/2)$. This completes the proof
of \eqref{sinmax}.

For the proof of \eqref{sinold}, let $$j(x)=\log\left(\frac{x}{\sin(x)}-\frac{2}{3}\frac{2}{1+\cos(x)}\right),$$
$x\in(0,\pi/2)$.
One has
\begin{eqnarray*}
j'(x)&=&\frac{1-x\cot(x)}{x}-\frac{2}{3}\frac{\sin(x)}{1+\cos(x)}\\
&=&\frac{1}{x}-\cot(x)-\frac{2}{3}\frac{1-\cos(x)}{\sin(x)}=j_1(x),
\end{eqnarray*}
which is negative, because the inequality  $j_1(x)<0$ can be written as $\sin(x)/x<(2+\cos(x))/3$,
which is so-called Cusa-Huygens inequality \cite{mit}. Thus,
$j(x)$ is strictly decreasing, and
$$\lim_{x->0}j(x)=0>j(x)>\log(\pi)-(5/3)\log(2)=\lim_{x\to \pi/2}j(\pi/2)\approx -0.010515.$$
By these inequalities we get
$$\cos(x/2)^{4/3}<\frac{\sin(x)}{x}<\exp((5/3)\log(2)-\log(\pi))\cos(x/2)^{4/3}.$$
The proof of the first and second inequality is completed, and the proof of the third inequality is trivial.
\end{proof}

The inequality \eqref{sinmax} improves the following one
\begin{equation}\label{sandor1702}
1-2\frac{\pi-2}{\pi^2}x<\frac{\sin(x)}{x},\quad 0<x<\pi,
\end{equation}
which was proved in \cite{sandor17} as an application of the concavity
of $\sin(x)/x$. Indeed, the inequality
$$1-2\frac{\pi-2}{\pi^2}x<\frac{2}{\pi}+\frac{\pi-2}{\pi}\cos(x),$$
is equivalent to
\begin{equation}\label{kober1702}
\cos(x)>1-\frac{2}{\pi}x,\quad 0<x<\frac{\pi}{2},
\end{equation}
which is Kober's inequality, see \cite{mit,sandor17b}.



%

\vspace{.3cm}
\noindent{\bf Proof of Theorem \ref{thm1702}.} \rm For $x\in(0,\pi/2)$, let
$$f(x)=\frac{x}{\sin(x)}\frac{x(\pi-x)}{(2x-\pi)^2}-\frac{4x^2}{(2x-\pi)^2}.$$
We get,
$$f'(x)=-\frac{16\pi^2\cos(x)}{(2x-\pi)^3\sin(x)^2}g(x),$$
where
$$g(x)=\tan(x)-\frac{\sin(x)^2}{\cos(x)}-x.$$
One has,
$$g'(x)=\frac{\sin(x)^2}{\cos(x)}(\sin(x)^2-\sin(x)-2)<0,$$
as $0<\sin(x)<1$, and $g(x)=0$. Thus, $g<0$, and in result $f'(x)<0$, this implies that
$f$ is strictly decreasing. By l'H\^opital rule we get
$$\lim_{x\to \pi/2}f(x)=\frac{\pi^2-8}{2}<f(x)<\lim_{x\to 0}f(x)=\frac{16}{\pi}-4,$$
this implies the proof. $\hfill\square$

\vspace{.5cm}
Replacing $x$ by $\pi/2-x$ in the inequalities of Theorem \ref{thm1702}, we get the following corollary as a result.

\begin{corollary}\label{cor1702} For $x\in(0,\pi/2)$, we have
$$\frac{16(\pi-x)}{4\pi^2+\alpha (2x-\pi)^2}<\frac{\sin(x)}{x}<\frac{16(\pi-x)}{4\pi^2+\beta (2x-\pi)^2},$$
where $\alpha$ and $\beta$ are as in Theorem \ref{thm1702}.
\end{corollary}


\end{document}